# Embedded FPGA Developments in 130nm and 28nm CMOS for Machine Learning in Particle Detector Readout

### v2

@@ 1 Background
In the sciences, many data processing tasks have a requirement to perform sophisticated AI inference within resource-constrained systems, such as with low power or latency, motivating the incorporation of ML into microelectronics. This is enabled by “fast ML", wherein high-level synthesis of software algorithms enables the implementation of ML models in hardware. In this work the embedded FPGA (eFPGA) technology is studied as an option to dynamically reconfigure logic to enable ML within an ASIC. eFPGAs comprise a digital block of FPGA fabric that can be embedded into an ASIC design. In this way they can offer the best of both worlds between FPGA and ASIC processing, with lower power demand than an FPGA but with the ease of firmware-based reconfiguration. eFPGAs also mitigate issues of commercial FPGA obsolescence by providing a hardware structure that can be maintained without industrial support, making it easier to maintain computational systems over the long lifetime of a scientific experiment.
In high energy collider facilties like the Large Hadron Collider (LHC) at the CERN laboratory in Geneva, Switzerland, the ability to execute reconfigurable and fast logic is indispensable. The LHC collides bunches of protons at a frequency of 40 MHz, creating thousands of particles that pass through detectors surrounding the beam interaction point. These conditions create a requirement to handle a total data rate of hundreds of terabytes of data per second, with latencies as low as $\mathcal{O}$(ns). Furthermore, the high radiation doses and stringent spatial constraints associated to the collision environment motivate the development of custom ASICs and electronics to read out the detector channels at high speed with good fidelity.
@@ 2.4.2 ASIC Power Draw
To measure the power draw, the same firmware from the 16-bit counter test was used, and the clock frequency of the ASIC was varied. This clock generated by the KCU105’s FPGA is sourced by the ASIC’s SUGIO, AXI-Lite endpoints, and the eFPGA itself. At various clock frequencies ranging from 10 to 125 MHz, both the ASIC’s core rail (+1.2VDC) and I/O rail (+1.2VDC) current draws were measured to calculate the total ASIC power drawing. Plots of these measurements are shown in Figure 5. The core voltage power consumption was observed to increase linearly with the clock frequency. This linear relationship is influenced by the frequency dependence of the switching events in the circuit. The non-linear power consumption for the I/O voltage is due to the load of ASIC PCB carrier and KCU105’s FPGA, which causes an incomplete settling of the digital signals at high frequencies (while still working correctly), thus reducing the current consumption.
@@
Figure 5: Plot of 130nm ASIC core voltage power draw versus clock frequency (left), and plot of 130nm ASIC I/O voltage power draw versus clock frequency (right).
@@ 4.4.2 ASIC Power Draw
Figure 10: Plot of 28nm ASIC core voltage power draw versus clock frequency (left), and plot of 28nm ASIC I/O voltage power draw versus clock frequency (right).
@@ 5 Application for Machine Learning-based At-Source Processing
The “smart pixel” collaboration provides a simulated dataset of particles propagated through a futuristic pixel sensor, which can be utilized to study ML-based readout [24]. This dataset comprises about 620,000 fitted tracks originating from high-energy pions collected by the Compact Muon Solenoid (CMS) experiment and propagated through a futuristic pixel detector. This detector features sensors composed of a 21x13 pixel array with a 50 x 12.5 $\mu$m pitch, located at a radius of 30 mm from the beamline within a magnetic field of 3.8 T. Each track is represented as a sequence of eight deposited charge arrays in the (x, y) pixel dimensions, at time intervals of 200 ps. Figure 11 illustrates a diagram of a pixel sensor from this dataset, and an example signal from a charged particle track passing through it.
@@
The performance of the BDT for the pileup classification task is evaluated using the test dataset. The background rejection, defined as the fraction of pileup tracks (with truth $p_{T}<$ 2 GeV) that are correctly classified as background and thus rejected, is evaluated for a given signal efficiency, defined as the fraction of tracks with truth $p_{T}>2$ GeV that are correctly classified as signal. Correct classification is determined by threshold setting, which determines the value of the output BDT probability that defines a signal classification. Before quantization, a background rejection of 4.35% is achieved for a signal efficiency of 97.53% for classification threshold $=0.4922$. After synthesis with quantization using ap_fixed<28,19>, the performance of the model under varying classification thresholds is shown in Table 1.
@@
Table 1: Performance of the synthesized and quantized BDT model under different classification thresholds, defined by the BDT output probability.
@@
While the performance of this model is limited by the small logical capacity of the hardware, its ability to successfully perform the classification task and fit the resource limitations enabled a full configuration test on the 28nm eFPGA. High-level synthesis of the software algorithm was performed using Conifer [28], an open-source framework for generating firmware configuration files from software-based BDT algorithms. A full hardware synthesis of the pileup classification BDT was achieved, including quantization of split threshold values in BDT and pruning to accommodate the BDT within stringent resource constraints, as well as a hardware test comparing the hardware output and the expected output from the quantized software version of the model. This process involved synthesis from C to Verilog firmware and simulation of the hardware response. A single tree was synthesized into a single decision function module in RTL, featuring only 9 threshold parameters and 14 inputs. The entire module utilized 294 LUTs, fitting within the confines of the current 28nm eFPGA design. The compiled bitstream was then loaded into the eFPGA, and a test of the BDT evaluation was performed using the full smart pixel dataset of 620,000 tracks. The model output when run in hardware is 100% accurate with respect to the expected output from the quantized software version of the synthesized model, achieving an operational runtime of less than 25 ns in simulation.
These results provide a proof-of-concept for the use of ML on eFPGAs, confirming the ability to use the open-souce FABulous framework for eFPGA designs and opening the door to a variety of future applications in the sciences. However, further development is needed to enable eFPGA technology for readout in future collider detectors. The small logical capacities of the ASICs described here are not feasible for data processing at future HEP particle detectors, which need to read out $\mathcal{O}$(105) pixels with very good signal efficiency and better background rejection than is provided by the BDT described here. A next-generation eFPGA with a larger logical capacity could enable the study of higher-performance models and a variety of algorithms to fully exploit the configurability of the ASIC. Further studies into power consumption will also be needed to determine whether the eFPGA can be sufficiently efficient for the spatially constrained detector environment. Additionally, any readout ASIC in a collider experiment will need to be insensitive to radiation-induced issues such as single-event effects. The implementation of triple modular redundancy (TMR) in FABulous could open up a broad usage of eFPGAs in collider experiment readout scenarios.

### v4

@@ 1 Background
In the sciences, many data processing tasks have a requirement to perform sophisticated AI inference within resource-constrained systems, such as with low power or latency, motivating the incorporation of ML into microelectronics. This is enabled by “fast ML", wherein high-level synthesis of software algorithms enables the implementation of ML models in hardware. In this work the embedded FPGA (eFPGA) technology is studied as an option to dynamically reconfigure logic to enable ML within an ASIC. eFPGAs comprise a digital block of FPGA fabric that can be embedded into an ASIC design. In this way they can offer the best of both worlds between FPGA and ASIC processing, with lower power demand than an FPGA but with the ease of firmware-based reconfiguration. eFPGAs also mitigate issues of commercial FPGA obsolescence by providing a hardware structure that can be maintained without industrial support, making it easier for computational systems to be maintained over the long lifetime of a scientific experiment.
In high energy collider facilities like the Large Hadron Collider (LHC) at the CERN laboratory in Geneva, Switzerland, the ability to execute reconfigurable and fast logic is indispensable. The LHC collides bunches of protons at a frequency of 40 MHz, creating thousands of particles that pass through detectors surrounding the beam interaction point. These conditions create a requirement to handle a total data rate of hundreds of terabytes of data per second, with latencies as low as $\mathcal{O}$(ns). Furthermore, the high radiation doses and stringent spatial constraints associated to the collision environment motivate the development of custom ASICs and electronics to read out the detector channels at high speed with good fidelity.
@@ 2.4.2 ASIC Power Draw
To measure the power draw, the same firmware from the 16-bit counter test was used, and the clock frequency of the ASIC was varied. This clock generated by the KCU105’s FPGA is sourced by the ASIC’s SUGIO, AXI-Lite endpoints, and the eFPGA itself. At various clock frequencies ranging from 10 to 125 MHz, both the ASIC’s core rail (+1.2VDC) and I/O rail (+1.2VDC) current draws were measured to calculate the total ASIC power drawing. Plots of these measurements are shown in Figure 5. The core voltage power consumption was observed to increase linearly with the clock frequency. This linear relationship is influenced by the frequency dependence of the switching events in the circuit. The non-linear power consumption for the I/O voltage is due to the load of the ASIC PCB carrier and KCU105’s FPGA, which causes an incomplete settling of the digital signals at high frequencies (while still working correctly), thus reducing the current consumption.
@@
Figure 5: Plot of the 130nm ASIC core voltage power draw versus clock frequency (left), and plot of the 130nm ASIC I/O voltage power draw versus clock frequency (right).
@@ 4.4.2 ASIC Power Draw
Figure 10: Plot of the 28nm ASIC core voltage power draw versus clock frequency (left), and plot of the 28nm ASIC I/O voltage power draw versus clock frequency (right).
@@ 5 Application for Machine Learning-based At-Source Processing
The “smart pixel” collaboration provides a simulated dataset of particles propagated through a futuristic pixel sensor, which can be utilized to study ML-based readout [24]. This dataset comprises about 550,000 fitted tracks originating from high-energy pions collected by the Compact Muon Solenoid (CMS) experiment and propagated through a futuristic pixel detector. This detector features sensors composed of a 21x13 pixel array with a 50 x 12.5 $\mu$m pitch, located at a radius of 30 mm from the beamline within a magnetic field of 3.8 T. Each track is represented as a sequence of eight deposited charge arrays in the (x, y) pixel dimensions, at time intervals of 200 ps. Figure 11 illustrates a diagram of a pixel sensor from this dataset, and an example signal from a charged particle track passing through it.
@@
The performance of the BDT for the pileup classification task is evaluated using the test dataset. The background rejection, defined as the fraction of pileup tracks (with truth $p_{T}<$ 2 GeV) that are correctly classified as background and thus rejected, is evaluated for a given signal efficiency, defined as the fraction of tracks with truth $p_{T}>2$ GeV that are correctly classified as signal. The performance of the classifier is determined by setting a threshold for the BDT output above or below which a track is classified as signal or background, respectively. Before quantization, a background rejection of 4.35% is achieved for a signal efficiency of 97.53% for a classification threshold of $0.4922$. After synthesis with quantization using ap_fixed<28,19>, the performance of the model under varying classification thresholds is shown in Table 1. Quantization can lead to slight discrepancies in performance for a fixed threshold.
@@
Table 1: Performance of the synthesized and quantized BDT model under different classification thresholds, defined by the BDT output.
@@
While the performance of this model is limited by the small logical capacity of the hardware, its ability to successfully perform the classification task and fit the resource limitations enabled a full configuration test on the 28nm eFPGA. High-level synthesis of the software algorithm was performed using Conifer [28], an open-source framework for generating firmware configuration files from software-based BDT algorithms. A full hardware synthesis of the pileup classification BDT was achieved, including the quantization of split threshold values in BDT and pruning to accommodate the BDT within stringent resource constraints. In addition, the hardware output was successfully compared to the expected output from the quantized software version of the model. This process involved the synthesis from C to Verilog firmware and the simulation of the hardware response. A single tree was synthesized into a single decision function module in RTL, featuring only 9 threshold parameters and 14 inputs. The entire module utilized 294 LUTs, fitting within the confines of the current 28nm eFPGA design. The compiled bitstream was then loaded into the eFPGA, and a test of the BDT evaluation was performed using the full smart pixel dataset of approximately 550,000 tracks. The model output when run in hardware was found to be 100% accurate with respect to the expected output from the quantized software version of the synthesized model, with an operational runtime of less than 25 ns in simulation.
These results provide a proof-of-concept for the use of ML on eFPGAs, confirming the ability to use the open-souce FABulous framework for eFPGA designs and opening the door to a variety of future applications in the sciences. However, further development is needed to enable eFPGA technology for readout in future collider detectors. The small logical capacities of the ASICs described here are not feasible for data processing at future HEP particle detectors, which need to read out $\mathcal{O}$(105) pixels with very good signal efficiency and better background rejection than what is provided by the BDT described here. A next-generation eFPGA with a larger logical capacity could enable the study of higher-performance models and a variety of algorithms to fully exploit the configurability of the ASIC. Further studies into power consumption will also be needed to determine whether the eFPGA can be sufficiently efficient for the spatially constrained detector environment. Additionally, any readout ASIC in a collider experiment will need to be insensitive to radiation-induced issues such as single-event effects. The implementation of triple modular redundancy (TMR) in FABulous could open up a broad usage of eFPGAs in collider experiment readout scenarios.